OpenStudy (anonymous):

quadratic formula can someone show me how to do this problem please. 4X to the power of 2-10X=8X thank you.

hero (hero):

4x^2 - 10x = 8x 4x^2 - 10x - 8x = 0 4x^2 - 18x = 0 4x(x - 4.5) = 0 4x = 0 x - 4.5 = 0 x = 4.5

hero (hero):

x = 0, 4.5

OpenStudy (anonymous):

4X^2-10X=8X thank you!!

OpenStudy (anonymous):

how about this problem 9X^2-10=6

hero (hero):

9x^2 - 10 = 6 9x^2 = 6 + 10 9x^2 = 16 x^2 = 16/9 x = sqrt(16/9) x = -4/3 x = +4/3

OpenStudy (anonymous):

thx!

hero (hero):

yw